OpenStudy (ny,ny):

Kim rode her bike at a rate of 24 km/h to the post office and 16 km/h back home. If her trip to the post office and back took a total of 75 minutes how far does Kim live from the post office?

OpenStudy (whpalmer4):

Let's say x is the distance to the post office from Kim's home. Kim rides 2x km in 75 minutes. Half of the distance she rides at 24 km/h, and half she rides at 16 km/h. The time to ride the trip there is x (km) / 24 (km/h) = x/24 (h). The time to ride back is x (km) / 16 (km/h) = x/16 (h). We know that the two trips add up to 75 minutes. 75 minutes * 1 hour / 60 minutes = 5/4 hours. \[\frac{x}{24} + \frac{x}{16} = \frac{5}{4}\] Can you solve that for \(x\)?

OpenStudy (ny,ny):

not sure how to.

OpenStudy (stamp):

@NY,NY currently those fractional terms each have an uncommon denominator. Fractional terms can only be combined once they have like-denominators. What is your LCD, or Least Common Denominator?

OpenStudy (ny,ny):

2.

OpenStudy (stamp):

2 is a common multiple, but is not your least common denominator.

OpenStudy (ny,ny):

1?

OpenStudy (stamp):

Ok, forget trying to find a Least Common Denominator. What if we took the whole equation and multipled each term by 24? Try that.

OpenStudy (whpalmer4):

You can also multiply the entire equation by 48, which is the least common multiple of those numbers. You can find the least common multiple by just writing like this: 4 8 12 16 20 24 28 32 36 40 44 48 52 56 16 32 48 64 24 48 72 48 is the first number that appears in each of the lists of multiples. \[\frac{x}{24} + \frac{x}{16} = \frac{5}{4}\]Multiply by 48 gives \[\frac{48x}{24} + \frac{48x}{16} = \frac{240}{4}\]And dividing those out \[2x + 3x = 60\] Now can you solve it?

OpenStudy (anonymous):

24*((5/4)*hr*(1/2))*km/hr+16*((5/4)*hr*(1/2))*km/hr =25 km

OpenStudy (whpalmer4):

@timo86m how about not just posting answers, especially wrong answers :-)

OpenStudy (ny,ny):

@whpalmer4 5x=60 x=12

OpenStudy (whpalmer4):

RIght! So, we have an answer. Now is the important part: checking it! If she rides 12 km at 24 km/h, it takes 1/2 hour (12/24). In the other direction, she rides 12 km at 16 km/h, so it takes (12/16) = 3/4 hour. 1/2 + 3/4 = 2/4 + 3/4 = 5/4. Our answer checks out.

OpenStudy (whpalmer4):

Always, always, always check your answers for these problems...

OpenStudy (ny,ny):

right. thanks :)

OpenStudy (whpalmer4):

And it's worth developing a sense of what a correct answer might be, too. timo86m said 25 km was the answer. Now, does that make sense? She rides her bike at only 24 km/h when she's going fast, so the fast trip would take a little more than an hour. There's no way she's going to get home in a quarter of an hour riding slowly, is there?

OpenStudy (ny,ny):

nope.